[Hallc_running] [New Logentry] Hall C RC Update for Wed

brads at jlab.org brads at jlab.org
Wed Jul 3 09:20:08 EDT 2019

Logentry Text:
Things are moving pretty well in MCC.  Cryo had the damaged relief valve for the 4K cold box repaired yesterday and they will reinstall it this morning.  If that goes well, 2K pumpdown and cryomodule top-offs will begin this afternoon.  I gather our He inventory is OK.  There are a few punchlist items that OPs are working throughout the machine, but no major pending issues.

As far as I can tell, Hall C is basically ready to go (just have to cool the target at an appropriate time).  I'll be in touch with Dave M. about this.

We will plan to lock up the Hall by the end of Day shift today for power-supply hot checkout.  We'll return to Controlled Access when they are done their thing.

** If you are in charge of a Hall checklist, please walk through your items and make an HCLog entry that cc's Walter, Brad, and Dave G. **

** Wednesday (today) SWING, Thurs OWL shifts are cancelled.  Thursday DAY and SWING are on standby.  We'll know more after the 1:30pm MCC meeting today **


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3698308
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190703/0b00b7b7/attachment.html>

More information about the Hallc_running mailing list